Solve (3x-1)1/3=3


1
Expert's answer
2021-03-19T04:19:08-0400

Given:

"(3x-1)^{\frac{1}{3}}=3"


"((3x-1)^{\frac{1}{3}})^\blue{3}=3^\blue{3}"

"((3x-1)^{\red{\frac{1}{3}}})^\red{3}=27"


Use the following basic rule for exponentiation:

"(a^{b})^c=a^{b \cdot c}"


"(3x-1)^\red{3 \cdot\frac{1}{3}}=27"

"(3x-1)^1=3x-1=27"

"3x-1=27"

"3x-1+\blue1=27+\blue1"

"3x=28"

"x=\frac{28}{3}"


The solution of the given equation:

"x=\frac{28}{3}"